OpenStudy (mathhelp346):

Use graphing to solve: a. |x+2|=6 b.|x+2|>6

OpenStudy (anonymous):

\[|x+2|=6\implies x +2 = 6 \text { or } x+2 = -6\] so ' \[x=-4\] or \[x=-8\]

OpenStudy (anonymous):

no that is wrong \[x=4\] or \[x=-8\]

OpenStudy (mathhelp346):

no i need to graph it, but i dont know how

OpenStudy (anonymous):

the graph is just two numbers on the number line

OpenStudy (mathhelp346):

??

OpenStudy (mathhelp346):

no i mean like the V thing

OpenStudy (anonymous):



OpenStudy (anonymous):

you mean graph \[y=|x+2|\]??

OpenStudy (mathhelp346):

like this